TIME-RESOLVED FOURIER TRANSFORM INFRARED EMISSION SPECTROSCOPY OF LASER ABLA-TION PRODUCTS

\underline{K. KAWAGUCHI}, N. SANECHIKA, Y. NISHIMURA, R. FUJIMORI, Faculty of Science, Okayama University, Okayama 700-8530, Japan; T. N. OKA, Y. HIRAHARA, Graduate School of Environmental Studies, Chikusa, Nagoya, 464-8601, Japan; A. I. JAMAN, Saha Institute of Nuclear Physics, Bidhannagar, Kolkata -700 064, India; and S. CIVIS, J. Heyrovsky Institute of Physical Chemistry, Dolejskova 3, 18223 Prague 8, Czech Republic.

Time-resolved Fourier transform infrared spectroscopy^{a} was applied for observations of emission spectra from Nd:YLF laser ablation products. A high-repetition rate laser (Photonics Industry DM10-527, 2.5 kHz) with 1-2 mJ output power was focused on a rotating rod of metal and carbon mounted on a linear actuator. The infrared emission spectra from Fe, Cu, Zn, and Al atoms were observed in the 2.5-5 μm region with a time profile showing maximum emission intensity in 3-5 μsec after a laser shot. The time profile was explained by diffusion process. The observed emission spectrum from iron ablation in the 2500 cm^{−1} region agrees very well with solar absorption spectrum, where new lines have been detected in the present experiment in addition to the lines observed from a hollow cathode discharge^{b}. When O_{2} was added to the carbon ablation, emissions from vibrationally excited CO was observed with a vibrational temperature of ∼12300 K. Based on the observed time profiles, relaxation processes will be discussed.

aK. Kawaguchi, O. Baskakov, Y. Hosaki, Y. Hama, and C. Kugimiya, J. Chem. Phys. Letters 369, 293 (2003).

bG. Nave, S. Johansson, R. C. M. Learner, A. P. Thorne, and J. W. Brault, Astrophy. J. Suppl. Ser. 94, 221 (1994).